A firm has made a credit sales with the following credit terms: $2,000,000 order, 30-day credit...

  1. A firm has made a credit sales with the following credit terms:
  • $2,000,000 order, 30-day credit terms
  • VCR (0.30) = Variable cost ratio/$ of sales
  • EXP (0.05/CP) = Expenses for credit administration and collection/$ of sales
  • i (0.10/365) = Daily interest rate
  • CP (45 days) = Collection period for sale
  1. Calculate the variable cost of this sales
  2. Calculate the present value of this sales
  3. Calculate the NPV of this sales
Answer #1

Part (a)

Variable cost = 0.30 x Sales = 0.30 x $ 2,000,000 = $  600,000

Part (b)

PV of the sales = Sale Value / (1 + daily interest rate)credit days = 2,000,000 / (1 + 0.10 / 365)30 = $  1,983,631

Part (c)

NPV = PV of sales - variable cost - Expenses for credit administration and collection = 1,983,631 - 600,000 - 0.05 x 2,000,000 = $  1,283,631